What volume of 0.220 MM ethanol solution contains each of the following number of moles of ethanol?

a) 0.47 molmol ethanol

5) 1.20 molmol ethanol

c) 1.2×10−2 molmol ethanol
1
Expert's answer
2020-02-19T07:57:35-0500

Molarity"(M)=\frac{N_{mol}}{V_{sol}}"

where "N_{mol}" is number of moles of solute and "V_{sol}" is volume of solution.

"\implies V_{sol}=\frac{N_{mol}}{M}"

Given,The solution has "M=0.220 \ mol/L" .

"a)" "0.47" mol ethanol 

"N_{mol}=0.47"

"\implies V_{sol}=\frac{0.47}{0.220}=2.13\ L"

"b) \ 1.20" mol ethanol

"N_{mol}=1.20"

"\implies V_{sol}=\frac{1.20}{0.220}=5.45\ L"

"c)\ 1.2\times 10^{-2}" mol ethanol

"N_{mol}=1.2\times 10^{-2}"

"\implies V_{sol}=\frac{1.2\times10^{-2}}{0.220}=5.45\times 10^{-2}\ L=54.5\ ml"